# A Methodology for Incompleteness-Tolerant and Modular Gradual Semantics for Argumentative Statement Graphs

### v5

@@ 3.3 Theoretical Analysis
11: Rejected Premise $\checkmark$ $\checkmark$ $\checkmark$ $\times$ -- --
12: Accepted Premises $\times$ $\times$ $\checkmark$ $\times$ -- --
@@
Table 1: Structured ($\sigma_{T_{p}}$, $\sigma_{T_{m}}$, $\sigma_{\wedge_{D}}$ and $\sigma_{\wedge_{Q}}$) and abstract ($\sigma_{D}$ and $\sigma_{Q}$) GS’ satisfaction (denoted with $\checkmark$) or violation (denoted with $\times$) of the assessed properties, where incompatibility is denoted with --.
@@
We now assess whether instantiations of the DC semantics, as well as existing structured and abstract GS, satisfy the aforementioned properties. Table 1 summarises the results. [^9] Full propositions and proofs are given in Appendix B.
@@
First, we assess the T-norm semantics in §2, $\sigma_{T_{p}}$ and $\sigma_{T_{m}}$. Both GS satisfy identical properties across our collection, however, they perform more poorly when we consider the novel properties. Even stability and neutrality, two basic properties in the realm of abstract GS, are violated due to the fact that they do not consider the weight as a starting point, nor the strength as a recursive function, respectively. Attacked and supported premises are both satisfied, while only weakened, and not strengthened, premise is satisfied due to the asymmetry in the way the attackers and supporters are evaluated, which leads to mirroring being violated also. Similarly, rejected premises is satisfied, but accepted premises is violated due to the (potentially non-maximal) weight of the statement playing a role in the calculation. Given accepted premises’ fundamental role across for statements of any completeness, this represents a significant weakness.
@@
Next, we consider the DC semantics supported by either the DF-QuAD semantics ($\sigma_{\wedge_{D}}$) or the QEM semantics ($\sigma_{\wedge_{Q}}$). Both DC semantics only satisfy directionality of the existing properties. Examples 3 and 4 demonstrate what we believe is reasonable justification for this. For example, rewriting is not satisfied by either semantics since more evidence, even if the statements constituting it are not complete, will strengthen an statement, which seems to be reasonable particularly for our settings with partially-complete or incomplete statements. Moreover, provability is violated by design, since we wanted GS that do not necessarily reject a statement when its supporting evidence is missing, only doing so when the statement’s weight is also minimal, hence the semantics’ satisfaction of weak provability instead. Both GS align with stability in treating the weight as a “starting point”, and from there they satisfy attacked and supported premises, as well as weakened and strengthened premises, due to the recursive way they handle attackers and supporters. Interestingly, the DC semantics, when supported by the DF-QuAD semantics, satisfies rejected premises and accepted premises, but it does not when supported by the QEM semantics. This is due to the saturation effect that is seen at the extremities of the strength scale in the DF-QuAD semantics, giving rise to this intuitive behaviour for this specific type of SG (with conjunctions as premises). Finally, mirroring is satisfied by both GS.
@@
We now assess the two abstract GS, the DF-QuAD semantics ($\sigma_{D}$) and the QEM semantics ($\sigma_{Q}$), as defined in §2 along the compatible properties. Although both of the abstract GS satisfy all of these properties, they are incompatible with many others, showing the advantages of structured in considering statements’ logical structure.
@@
While both T-norm semantics perform well in the existing properties which require complete information, under incomplete information, i.e., with partially-complete or incomplete statements, where the novel properties are particularly powerful, there are clear weaknesses. This highlights how the DC semantics, and in particular that supported by DF-QuAD, offer clear advantages. Thus, in settings with varying completeness of information, our approach fills a gap in the literature.
@@
4 Related Work
Abstract GS have received significant attention in recent years, given their professed alignment with human reasoning Vesic et al. (2022); de Tarlé et al. (2022) and applicability to real-world settings Potyka (2021); Rago et al. (2021). This has given rise to a rich field of research, with a broad repertoire of GS Gonzalez et al. (2021); Yun and Vesic (2021); Wang and Shen (2024), in-depth analyses of GS’ behaviour Delobelle and Villata (2019); Oren et al. (2022); Yin et al. (2023); Kampik et al. (2024), and a number of open-source computational toolsets Potyka (2022); Alarcón and Kampik (2024). One issue in this area is that many abstract GS, given their recursive nature, do not satisfy the uniqueness property in assigning strengths in cyclic graphs Gabbay and Rodrigues (2015a); Anaissy et al. (2024). This must be considered if our acyclicity restriction is relaxed in future work.
With regards to other approaches to GS in structured argumentation, our approach differs from existing work as follows. Prakken Prakken (2024) gives a formal model of argument strength and deploys it within ASPIC+ but does not consider the support relation and thus cannot be studied with respect to many of our properties. Meanwhile, Heyninck et al. Heyninck et al. (2023) compare properties of GS with those concerning culpability when applied to abstract argumentation frameworks extracted from deductive argumentation, whereas we lift abstract GS to the structured level. Skiba et al. Skiba et al. (2023) assess a family of GS for ranking arguments in structured argumentation, specifically ABA. Both of these works ignore the support relation. Finally, Amgoud and Ben-Naim Amgoud and Ben-Naim (2015) give a method for ranking logic-based argumentation frameworks, but do not consider an explicit strength or a support relation.
5 Future Work
@@
Our methodology represents an early step forward in GS for structured argumentation, and opens various potentially fruitful directions for future work. First, we would like to broaden our analysis, both within our restricted form of SGs, exploring various other instantiations of our DC semantics, e.g., using other abstract GS such as the exponent-based restricted semantics Amgoud and Ben-Naim (2018), as well as extending to different forms of SGs with more complex logical premises. Specifically, our ability to deal with partially complete or incomplete statements may be related to the possibility of deriving assumptions in non-flat ABA Cyras et al. (2017). We would also like to consider whether other properties for abstract GS are desirable, e.g., open-mindedness Potyka (2019) (the satisfaction of which gives a clear advantage of $\sigma_{Q}$ over $\sigma_{D}$), attainability Cocarascu et al. (2019), whose suitability in our restricted SGs was not obvious due to the conjunctive premise. It would also be interesting to formalise the interplay between properties, e.g., between provability and stability, which are incompatible without restrictions. Further, it would be interesting to see whether our analysis could be applicable in the related but distinct subfield of probabilistic argumentation Kohlas (2003); Dung and Thang (2010); Hunter and Thimm (2014); Gabbay and Rodrigues (2015b); Fazzinga et al. (2018). Finally, we plan to explore how our methodology translates to real-world applications, e.g., in multi-agent model reconciliation Vasileiou et al. (2024) given GS’ capability for modelling human or machine reasoning Rago et al. (2023).
@@ Appendix A Additional Properties
In this section, we assess the GS against the properties for CSTs from Jedwabny et al. [2020]
@@ Appendix B Proofs
The proofs for Properties 1-3 can be found in Jedwabny et al. [2020].
@@
  • Property 11: Rejected Premise. Let us prove by contradiction. If we let $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{1})>0$, we can see from Definition 6 that $\exists T_{+}\in\mathcal{T}^{\mathcal{G}}(\alpha_{1})$ where $\mathcal{O}(T_{+})>0$. However, since $\alpha_{2}\in\mathcal{A}(\alpha_{1})$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{2})=1$, we can see from Definition 6 that $\exists T_{-}\in\mathcal{T}^{\mathcal{G}}(\alpha_{2})$ where $\mathcal{I}(T_{-})=1$ and exists $\alpha_{3}\in T_{-}$ such that $\tau(\alpha_{3})=1$. Thus, any $T_{+}$ must necessarily be attacked by $T_{-}$ and thus $\mathcal{O}(T_{+})=0$ and we have the contradiction.
  • Property 12: Accepted Premises (Counterexample). Let $\mathcal{X}=\{\alpha_{1},\alpha_{2}\}$, where $\alpha_{1}=\langle a,b\rangle$ and $\alpha_{2}=\langle\top,a\rangle$, and $\tau(\alpha_{1})=0.8$ and $\tau(\alpha_{2})=1$. Then, by Definition 6, we have that $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{2})=1$, but $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{1})=0.8$.
@@
The proofs for Properties 1-3 are analogous to those for $\sigma_{T_{p}}$ in Jedwabny et al. [2020].
@@
  • Property 11: Rejected Premise. Let us prove by contradiction. If we let $\sigma^{\mathcal{G}}_{T_{m}}(\alpha_{1})>0$, we can see from Definition 6 that $\exists T_{+}\in\mathcal{T}^{\mathcal{G}}(\alpha_{1})$ where $\mathcal{O}(T_{+})>0$. However, since $\alpha_{2}\in\mathcal{A}(\alpha_{1})$ and $\sigma^{\mathcal{G}}_{T_{m}}(\alpha_{2})=1$, we can see from Definition 6 that $\exists T_{-}\in\mathcal{T}^{\mathcal{G}}(\alpha_{2})$ where $\mathcal{I}(T_{-})=1$ and exists $\alpha_{3}\in T_{-}$ such that $\tau(\alpha_{3})=1$. Thus, any $T_{+}$ must necessarily be attacked by $T_{-}$ and thus $\mathcal{O}(T_{+})=0$ and we have the contradiction.
  • Property 15: Accepted Premises (Counterexample). Let $\mathcal{X}=\{\alpha_{1},\alpha_{2}\}$, where $\alpha_{1}=\langle a,b\rangle$ and $\alpha_{2}=\langle\top,a\rangle$, and $\tau(\alpha_{1})=0.8$ and $\tau(\alpha_{2})=1$. Then, by Definition 6, we have that $\sigma^{\mathcal{G}}_{T_{m}}(\alpha_{2})=1$, but $\sigma^{\mathcal{G}}_{T_{m}}(\alpha_{1})=0.8$.
@@
  • Property 4: Weak Provability. Let us prove by contradiction. If $\exists\alpha_{1}\in\mathcal{X}$ where $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{1})\neq 0$, then, by Definition 12, $\forall x\in Prem(\alpha_{1})$, $\sigma^{\mathcal{G}}_{D}(x)>0$. However, since, by the same definition, $\forall x\in Prem(\alpha_{1})$, $\tau(x)=\sqrt[n]{\tau(\alpha_{1})}$, where $n=|Prem(\alpha_{1})|$ and $\tau(\alpha_{1})=0$, we know that for any $n$, $\tau(x)=0$. Then, since $\sigma_{D}$ satisfies balance [Baroni et al., 2019, Proposition 37], for $\sigma^{\mathcal{G}}_{D}(x)>0$ to hold it must be the case that $\exists\langle\Phi_{2},\Psi_{2}\rangle\in\mathcal{S}(\alpha_{1})$ s.t. $x\equiv\Psi_{2}$, by Definition 1, and thus we have the contradiction.
@@
  • Property 11: Rejected Premise. It can be seen from Definition 7 that $\Sigma(\sigma^{\mathcal{G}}_{D}(\mathcal{A}(x)))=1$ while $\Sigma(\sigma^{\mathcal{G}}_{D}(\mathcal{S}(x)))=0$. It follows from the same definition that, regardless of $\tau(x)$, $\sigma^{\mathcal{G}}_{D}(x)=0$. Then, by Definition 12, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{1})=0$.
  • Property 12: Accepted Premises. It can be seen from Definition 7 that $\forall x\in Prem(\alpha_{1})$, $\Sigma(\sigma^{\mathcal{G}}_{D}(\mathcal{A}(x)))=0$ while $\Sigma(\sigma^{\mathcal{G}}_{D}(\mathcal{S}(x)))=1$. It follows from the same definition that, regardless of $\tau(x)$, $\sigma^{\mathcal{G}}_{D}(x)=1$. Then, by Definition 12, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{1})=1$.
@@
  • Property 4: Weak Provability. Let us prove by contradiction. If $\exists\alpha_{1}\in\mathcal{X}$ where $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{1})\neq 0$, then, by Definition 12, $\forall x\in Prem(\alpha_{1})$, $\sigma^{\mathcal{G}}_{Q}(x)>0$. However, since, by the same definition, $\forall x\in Prem(\alpha_{1})$, $\tau(x)=\sqrt[n]{\tau(\alpha_{1})}$, where $n=|Prem(\alpha_{1})|$ and $\tau(\alpha_{1})=0$, we know that for any $n$, $\tau(x)=0$. Then, since $\sigma_{Q}$ satisfies weakening and strengthening [Potyka, 2018, Propositions 13 and 14], for $\sigma^{\mathcal{G}}_{Q}(x)>0$ to hold it must be the case that $\exists\langle\Phi_{2},\Psi_{2}\rangle\in\mathcal{S}(\alpha_{1})$ s.t. $x\equiv\Psi_{2}$, by Definition 1, and thus we have the contradiction.
@@
  • Property 11: Rejected Premise (Counterexample). Let $\mathcal{X}=\{\alpha_{1},\alpha_{2}\}$ where $\alpha_{1}=\langle b,a\rangle$ and $\alpha_{2}=\langle\top,\neg b\rangle$. By Definition 1, we have that $\mathcal{A}=\{(\alpha_{2},\alpha_{1})\}$ and $\mathcal{S}=\emptyset$. Then let $\tau(\alpha_{1})=0.5$ and $\tau(\alpha_{2})=1$. By Definitions 12, 11 and 8, this gives $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{2})=1$ and $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{1})=0.25$.
  • Property 12: Accepted Premises (Counterexample). Let $\mathcal{X}=\{\alpha_{1},\alpha_{2}\}$ where $\alpha_{1}=\langle b,a\rangle$ and $\alpha_{2}=\langle\top,b\rangle$. By Definition 1, we have that $\mathcal{A}=\emptyset$ and $\mathcal{S}=\{(\alpha_{2},\alpha_{1})\}$. Then let $\tau(\alpha_{1})=0.5$ and $\tau(\alpha_{2})=1$. By Definitions 12, 11 and 8, this gives $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{2})=1$ and $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{1})=0.75$.
@@
  • The proofs for Properties 14-17 can be found in Jedwabny et al. [2020].
@@
  • The proofs for Properties 14-17 are analogous to those for $\sigma_{T_{p}}$ in Jedwabny et al. [2020].

### v6

@@ 4.3 Theoretical Analysis
11: Bottom-Strength Premise $\checkmark$ $\checkmark$ $\checkmark$ $\times$ -- --
12: Top-Strength Premises $\times$ $\times$ $\checkmark$ $\times$ -- --
@@
Table 1: Structured ($\sigma_{T_{p}}$, $\sigma_{T_{m}}$, $\sigma_{\wedge_{D}}$ and $\sigma_{\wedge_{Q}}$) and abstract ($\sigma_{D}$ and $\sigma_{Q}$) GS’ satisfaction (denoted with $\checkmark$) or violation (denoted with $\times$) of the assessed properties, where incompatibility of a GS with a property is denoted with --.
@@
We now assess whether instantiations of the DC semantics, as well as existing structured and abstract GS, satisfy the properties from §4.2. Table 1 summarises the results.
First, we assess the T-norm semantics in §3, $\sigma_{T_{p}}$ and $\sigma_{T_{m}}$.
Proposition 1.
@@
Both GS satisfy identical properties across our collection, however, they perform more poorly when we consider the novel properties. Even stability and neutrality, two basic properties in the realm of abstract GS, are violated due to the fact that they do not consider the weight as a starting point, nor the strength as a recursive function, respectively. Attacked and supported premises are both satisfied, while only weakened, and not strengthened, premise is satisfied due to the asymmetry in the way the attackers and supporters are evaluated, which leads to mirroring being violated also. Similarly, bottom-strength premises is satisfied, but top-strength premises is violated due to the (potentially non-maximal) weight of the statement playing a role in the calculation. Given top-strength premises’ fundamental role across statements of any completeness, this represents a significant weakness.
Next, we consider the DC semantics, supported by either the DF-QuAD semantics ($\sigma_{\wedge_{D}}$) or the QEM semantics ($\sigma_{\wedge_{Q}}$), as defined in §4.1.
Proposition 2.
Both DC semantics only satisfy directionality of the existing properties. The violation of the other existing properties (rewriting and provability) can be justified by their incompatibility with the novel properties which we introduce.
Proposition 3.
Rewriting, stability and top-strength premises are incompatible.
Proposition 4.
Provability and stability are incompatible.
@@
This means that a user selecting a semantics must choose between its satisfation of one set of properties or another, so our approach represents a viable alternative to the existing work, particularly in cases with incomplete information. Examples 3 and 4 demonstrate what we believe is further justification for this. For example, rewriting is not satisfied by either semantics since more evidence, even if the statements constituting it are not complete, will strengthen an statement, which seems to be reasonable particularly for our settings with partially-complete or incomplete statements. Moreover, provability is violated by design, since we wanted GS that do not necessarily reject a statement when its supporting evidence is missing, only doing so when the statement’s weight is also minimal, hence the semantics’ satisfaction of weak provability instead. Both GS align with stability in treating the weight as a “starting point”, and from there they satisfy attacked and supported premises, as well as weakened and strengthened premises, due to the recursive way they handle attackers and supporters. Interestingly, the DC semantics, when supported by the DF-QuAD semantics, satisfies bottom-strength premise and top-strength premises, but it does not when supported by the QEM semantics. This is due to the saturation effect that is seen at the extremities of the strength scale in the DF-QuAD semantics, giving rise to this intuitive behaviour for this specific type of SG (with conjunctions as premises). Finally, mirroring is satisfied by both GS.
We now assess the two abstract GS, the DF-QuAD semantics ($\sigma_{D}$) and the QEM semantics ($\sigma_{Q}$), as defined in §3, along the compatible properties.
Proposition 5.
@@
Although both of the abstract GS satisfy all of the properties with which they are compatible, they are incompatible with many others, showing the advantages of structured GS in considering statements’ logical structure.
@@
While both T-norm semantics perform well in the existing properties which require complete information, under incomplete information, i.e., with partially-complete or incomplete statements, where the novel properties are particularly powerful, there are clear weaknesses. This highlights how the DC semantics, and in particular that supported by DF-QuAD, offer clear advantages, especially given that all of the properties cannot be satisfied by a given semantics. Thus, in settings with varying completeness of information, our approach fills a gap in the literature.
@@
5 Conclusions and Future Work
In this paper, we introduced a novel methodology for obtaining GS in SGs that naturally accommodate incomplete information. Our modular methodology dialectically evaluates the literals in a statement’s premise before aggregating these evaluations based on the statement’s construction. This separation allows our semantics to handle incomplete information by leveraging any existing GS for QBAFs to evaluate the available evidence without requiring complete support. We then demonstrated how our DC semantics can effectively leverage abstract GS for structured argumentation. Furthermore, we discussed existing and novel properties for GS in SGs, revealing some incompatibilities between existing properties, e.g., rewriting and provability, and our novel properties, e.g., stability and top-strength premises, showing that users must select between them given a particular contextual setting, e.g., based on the completeness notions we define. Our theoretical analysis demonstrated that the DC semantics supported by the DF-QuAD abstract GS satisfies all of our novel properties, while existing structured GS struggle with cases without complete information.
Note that while this paper establishes the foundations of our novel approach using a toy example for pedagogical reasons, we see significant potential for real-world applications, particularly in the analysis of debates mapped onto SGs through argument mining (?), possibly enhanced by large language models (??). It is well known that mining structured arguments is a complex problem, not least due to the presence of enthymemes causing incompleteness (??). A methodology leveraging GS that can handle incomplete information could alleviate some of the burden that structured argumentation’s requirements for completeness place on these methods.
@@ 5 Conclusions and Future Work
Our methodology represents an early step forward in GS for structured argumentation, and opens various potentially fruitful directions for future work. First, we would like to broaden our analysis, both within our restricted form of SGs, exploring various other instantiations of our DC semantics, e.g., using other abstract GS such as the exponent-based restricted semantics (?), as well as extending to different forms of SGs with more complex logical premises. Specifically, our ability to deal with partially-complete or incomplete statements may be related to the possibility of deriving assumptions in non-flat ABA (?). We would also like to consider whether other properties for abstract GS are desirable, e.g., open-mindedness (?) (the satisfaction of which gives a clear advantage of $\sigma_{Q}$ over $\sigma_{D}$), attainability (?), whose suitability in our restricted SGs was not obvious due to the conjunctive premise. It would also be interesting to formalise the interplay between properties, e.g., between provability and stability, which are incompatible without restrictions. Further, it would be interesting to see whether our analysis could be applicable to probabilistic argumentation (?????). Finally, we plan to explore how our methodology translates to real-world applications, e.g., in multi-agent model reconciliation (?) given GS’ capability for modelling human or machine reasoning (?).
@@ Appendix A Additional Properties
In this section, we assess the GS against the properties for CSTs from (?)
@@ Appendix B Proofs
The proofs for Properties 1-3 can be found in (?).
@@
  • Property 11: Bottom-Strength Premise. Let us prove by contradiction. If we let $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{1})>0$, we can see from Definition 6 that $\exists T_{+}\in\mathcal{T}^{\mathcal{G}}(\alpha_{1})$ where $\mathcal{O}(T_{+})>0$. However, since $\alpha_{2}\in\mathcal{A}(\alpha_{1})$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{2})=1$, we can see from Definition 6 that $\exists T_{-}\in\mathcal{T}^{\mathcal{G}}(\alpha_{2})$ where $\mathcal{I}(T_{-})=1$ and exists $\alpha_{3}\in T_{-}$ such that $\tau(\alpha_{3})=1$. Thus, any $T_{+}$ must necessarily be attacked by $T_{-}$ and thus $\mathcal{O}(T_{+})=0$ and we have the contradiction.
  • Property 12: Top-Strength Premises (Counterexample). Let $\mathcal{X}=\{\alpha_{1},\alpha_{2}\}$, where $\alpha_{1}=\langle a,b\rangle$ and $\alpha_{2}=\langle\top,a\rangle$, and $\tau(\alpha_{1})=0.8$ and $\tau(\alpha_{2})=1$. Then, by Definition 6, we have that $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{2})=1$, but $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{1})=0.8$.
@@
The proofs for Properties 1-3 are analogous to those for $\sigma_{T_{p}}$ in (?).
@@
  • Property 11: Bottom-Strength Premise. Let us prove by contradiction. If we let $\sigma^{\mathcal{G}}_{T_{m}}(\alpha_{1})>0$, we can see from Definition 6 that $\exists T_{+}\in\mathcal{T}^{\mathcal{G}}(\alpha_{1})$ where $\mathcal{O}(T_{+})>0$. However, since $\alpha_{2}\in\mathcal{A}(\alpha_{1})$ and $\sigma^{\mathcal{G}}_{T_{m}}(\alpha_{2})=1$, we can see from Definition 6 that $\exists T_{-}\in\mathcal{T}^{\mathcal{G}}(\alpha_{2})$ where $\mathcal{I}(T_{-})=1$ and exists $\alpha_{3}\in T_{-}$ such that $\tau(\alpha_{3})=1$. Thus, any $T_{+}$ must necessarily be attacked by $T_{-}$ and thus $\mathcal{O}(T_{+})=0$ and we have the contradiction.
  • Property 15: Top-Strength Premises (Counterexample). Let $\mathcal{X}=\{\alpha_{1},\alpha_{2}\}$, where $\alpha_{1}=\langle a,b\rangle$ and $\alpha_{2}=\langle\top,a\rangle$, and $\tau(\alpha_{1})=0.8$ and $\tau(\alpha_{2})=1$. Then, by Definition 6, we have that $\sigma^{\mathcal{G}}_{T_{m}}(\alpha_{2})=1$, but $\sigma^{\mathcal{G}}_{T_{m}}(\alpha_{1})=0.8$.
@@
  • Property 4: Weak Provability. Let us prove by contradiction. If $\exists\alpha_{1}\in\mathcal{X}$ where $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{1})\neq 0$, then, by Definition 12, $\forall x\in Prem(\alpha_{1})$, $\sigma^{\mathcal{G}}_{D}(x)>0$. However, since, by the same definition, $\forall x\in Prem(\alpha_{1})$, $\tau(x)=\sqrt[n]{\tau(\alpha_{1})}$, where $n=|Prem(\alpha_{1})|$ and $\tau(\alpha_{1})=0$, we know that for any $n$, $\tau(x)=0$. Then, since $\sigma_{D}$ satisfies balance (?, Proposition 37), for $\sigma^{\mathcal{G}}_{D}(x)>0$ to hold it must be the case that $\exists\langle\Phi_{2},\Psi_{2}\rangle\in\mathcal{S}(\alpha_{1})$ s.t. $x\equiv\Psi_{2}$, by Definition 1, and thus we have the contradiction.
@@
  • Property 11: Bottom-Strength Premise. It can be seen from Definition 7 that $\Sigma(\sigma^{\mathcal{G}}_{D}(\mathcal{A}(x)))=1$ while $\Sigma(\sigma^{\mathcal{G}}_{D}(\mathcal{S}(x)))=0$. It follows from the same definition that, regardless of $\tau(x)$, $\sigma^{\mathcal{G}}_{D}(x)=0$. Then, by Definition 12, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{1})=0$.
  • Property 12: Top-Strength Premises. It can be seen from Definition 7 that $\forall x\in Prem(\alpha_{1})$, $\Sigma(\sigma^{\mathcal{G}}_{D}(\mathcal{A}(x)))=0$ while $\Sigma(\sigma^{\mathcal{G}}_{D}(\mathcal{S}(x)))=1$. It follows from the same definition that, regardless of $\tau(x)$, $\sigma^{\mathcal{G}}_{D}(x)=1$. Then, by Definition 12, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{1})=1$.
@@
  • Property 4: Weak Provability. Let us prove by contradiction. If $\exists\alpha_{1}\in\mathcal{X}$ where $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{1})\neq 0$, then, by Definition 12, $\forall x\in Prem(\alpha_{1})$, $\sigma^{\mathcal{G}}_{Q}(x)>0$. However, since, by the same definition, $\forall x\in Prem(\alpha_{1})$, $\tau(x)=\sqrt[n]{\tau(\alpha_{1})}$, where $n=|Prem(\alpha_{1})|$ and $\tau(\alpha_{1})=0$, we know that for any $n$, $\tau(x)=0$. Then, since $\sigma_{Q}$ satisfies weakening and strengthening (?, Propositions 13 and 14), for $\sigma^{\mathcal{G}}_{Q}(x)>0$ to hold it must be the case that $\exists\langle\Phi_{2},\Psi_{2}\rangle\in\mathcal{S}(\alpha_{1})$ s.t. $x\equiv\Psi_{2}$, by Definition 1, and thus we have the contradiction.
@@
  • Property 11: Bottom-Strength Premise (Counterexample). Let $\mathcal{X}=\{\alpha_{1},\alpha_{2}\}$ where $\alpha_{1}=\langle b,a\rangle$ and $\alpha_{2}=\langle\top,\neg b\rangle$. By Definition 1, we have that $\mathcal{A}=\{(\alpha_{2},\alpha_{1})\}$ and $\mathcal{S}=\emptyset$. Then let $\tau(\alpha_{1})=0.5$ and $\tau(\alpha_{2})=1$. By Definitions 12, 11 and 8, this gives $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{2})=1$ and $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{1})=0.25$.
  • Property 12: Top-Strength Premises (Counterexample). Let $\mathcal{X}=\{\alpha_{1},\alpha_{2}\}$ where $\alpha_{1}=\langle b,a\rangle$ and $\alpha_{2}=\langle\top,b\rangle$. By Definition 1, we have that $\mathcal{A}=\emptyset$ and $\mathcal{S}=\{(\alpha_{2},\alpha_{1})\}$. Then let $\tau(\alpha_{1})=0.5$ and $\tau(\alpha_{2})=1$. By Definitions 12, 11 and 8, this gives $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{2})=1$ and $\sigma^{\mathcal{G}}_{\wedge_{Q}}(\alpha_{1})=0.75$.
@@
Proposition 3.
Properties 2, 5 and 12 are incompatible.
Proof.
Let us prove that Properties 2, 5 and 12 are incompatible by contradiction. Let us assume the existence of a semantics $\sigma_{a}$ which satisfies all three properties. Consider $\mathcal{X}=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ where $\alpha_{1}=\langle b,a\rangle$, $\alpha_{2}=\langle b,c\rangle$ and $\alpha_{3}=\langle c,a\rangle$. By Definition 1, we have that $\mathcal{A}=\emptyset$ and $\mathcal{S}=\{(\alpha_{2},\alpha_{3})\}$. Then let $\tau(\alpha_{1})=\tau(\alpha_{3})=0.5$ and $\tau(\alpha_{2})=1$. Then, Property 5 requires that $\sigma^{\mathcal{G}}_{a}(\alpha_{1})=0.5$, Property 2 requires that $\sigma^{\mathcal{G}}_{a}(\alpha_{1})=\sigma^{\mathcal{G}}_{a}(\alpha_{2})$, and Property 12 requires that $\sigma^{\mathcal{G}}_{a}(\alpha_{1})=0.5$, and so we have the contradiction.
Proposition 4.
Properties 3 and 5 are incompatible.
Proof.
Let us prove that Properties 3 and 5 are incompatible by contradiction. Let us assume the existence of a semantics $\sigma_{a}$ which satisfies both properties. Consider some $\alpha_{1}\in\mathcal{X}$ such that $\mathcal{A}(\alpha_{1})=\mathcal{A}(\alpha_{1})=\emptyset$ and $\tau(\alpha_{1})\neq 0$. Property 3 requires that $\sigma_{a}^{\mathcal{G}}(\alpha_{1})\neq 0=\tau(\alpha_{1})$, but Property 3 requires that $\sigma_{a}^{\mathcal{G}}(\alpha_{1})=0$, and we have the contradiction.
@@
  • The proofs for Properties 14-17 can be found in (?).
@@
  • The proofs for Properties 14-17 are analogous to those for $\sigma_{T_{p}}$ in (?).